1. Part B, C, D & E

1. Part B, C, D & E
Answer #1

b) b0 = 0.388+0.537*0.307 = 55.29

c) 55.29 % of the Bike riders to wear helmets

d) there is decrease of 0.537

e) For 40% of Children have free lunch

Estimated , % of Wearing helmts = 55.29-0.537*40 = 33.81%

Residual = 40-33.81 = 6.19